The area between the curve Y^2= 2x+5, the X -axis and the ordinates at X=1 and. X=3 is rotated about the X-axis. Calculate the volume of the solid generated.
1
Expert's answer
2019-09-20T11:34:27-0400

"V=\pi\int_1^3y^2dx=\pi\int_1^3(2x+5)dx=\pi(x^2+5x)|_1^3=18\pi."